40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 675     Entry time: Tue Jul 15 12:44:08 2008
Author: John 
Type: Summary 
Category: Locking 
Subject: DRFPMI with DC readout 
Rob, John

Last night, despite suspect alignment, we were again able to reduce the CARM offset to zero using
the RF signal.We were also able to transfer to dc readout taking calibrated spectra in both states.
DC readout shows a marked improvement over RF above ~1kHz but introduces some noise around 100Hz.
Broadband sensitivity appears to be more than ten times worse than previously. The calibration
being used remains to be confirmed.

Engaging the ETMY dewhitening caused lock to be lost. We'll check this today. The OMC alignment loops
may also need some attention.

We looked at REFL_166 as a candidate for CARM, at present POX still looks better.

The DARM filters were modified to reduce excess noise around 3Hz. Updating filter coefficients does
not cause loss of lock.
ELOG V3.1.3-